OpenStudy (anonymous):

I need to determine the convergence/divergence of this infinite series: \[\sum_{n=1}^{\infty} n! \div 2^{n}\]

OpenStudy (anonymous):

That diverges. Wildly, as it turns out.

OpenStudy (anonymous):

Yeah, my prof says it is diverging, but i need to know how it is derived.

OpenStudy (anonymous):

You can use any number of tests. Ratio test would probably be easiest: \[ \lim_{n -> \infty} \left| \frac{\frac{(n+1)!}{2^{n+1}}}{\frac{n!}{2^n}}\right| \] \[= \lim_{n \rightarrow \infty} \left| \frac{n+1}{2} \right| \] as n approaches infinity, that limit diverges, so the series diverges.

OpenStudy (anonymous):

Wow, that's cleared some things! Thanks a lot!